OpenStudy (anonymous):

A cyclist bikes at a constant speed for 21 miles . He then returns home at the same speed but takes a different route. His return trip takes one hour longer and is 26 miles. find the speed.

OpenStudy (anonymous):

(21/x)=(26/(x+1)) Solve for X

OpenStudy (anonymous):

I don't agree with this... Well, the equation is correct, but when you solve for x you are finding the time not the speed.

OpenStudy (anonymous):

speed = distance/time since the speed is constant, you set the speeds the same 21/x is the speed there and 26/(x+1) is the speed returning

OpenStudy (anonymous):

When you solve for x you get 21x + 21 = 26x 21 = 5x x = 4.2 this is time in hours. Substitute back in r = 21/x and r = 26/(t + 1) so rate = 21/4.2 = 5 mph and 26/(4.2+1) = 5 mph

OpenStudy (anonymous):

I agree you set the speed equal to each other, but the X represents the time in the equation not the rate. d = rt 21 = rt and 26 = r(t+1) solve for r r = 21/t and r = 26/(t+1) Set them equal, but you are solving for t which is time.....

OpenStudy (anonymous):

ahh...yea, my bad. You right